OpenStudy (anonymous):

Which of the following is not equivalent to the rational expression when x not equal to 7 ? x^2-49/x-7 A. x-7 B. x+7/x-7 C. 1/x+7 D. x+7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

a rather strangely worded question, which you can interpret as "factor and cancel" does that help?

OpenStudy (anonymous):

not really because I suck at math honestly and hi haha

OpenStudy (misty1212):

lol ok lets do it then factor the top, which is \(x^2-49\) the "difference of two squares" \[a^2-b^2=(a+b)(a-b)\] always, so \[x^2-49=(x+7)(x-7)\]

OpenStudy (anonymous):

which of the following is equal to the rational expression when x not equal to 7............ I TYPED THE QUESTION WRONG!!!

OpenStudy (misty1212):

i knew that since 3 of them are NOT and only one of them IS

OpenStudy (anonymous):

yes... sorry

OpenStudy (misty1212):

since the numerator is \((x+7)(x-7)\) the entire fraction is \[\frac{(x+7)(x-7)}{x-7}\] now cancel

OpenStudy (anonymous):

x+7?

OpenStudy (misty1212):

you know what i mean by that?

OpenStudy (misty1212):

yeah you get \(x+7\)

OpenStudy (anonymous):

thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

:)